Vocaloid:

Calc III: Standard parametrization for the **tangent** to the line r(t) = 2sin(t) i + (t^2 - cos(theta)) j + 2(e^(3t)) k, at t = 0

Vocaloid:

As a courtesy please do not respond unless you are making a serious, informed attempt to solve the problem. Thank you for being considerate of others.

Vocaloid:

what I have so far: r'(t) = 2cos(t) i + (2t - sin(theta))j + 6e^(3t) r(0) = (0,-1,2) r'(t) = r(0) + t*r'(t) ?

sillybilly123:

+ sin t

sillybilly123:

in the j bit